OpenStudy (anonymous):

Given the function g(x) = −4x + 5, find g(−4). −11 21 2.25 11

OpenStudy (igloo3000):

2?

OpenStudy (igloo3000):

i don't know how to solve?

OpenStudy (anonymous):

Substitute -4 for x. You get -4(-4) + 5 = 16 + 5 = 21

OpenStudy (anonymous):

Damn....no medals huh? :)

OpenStudy (anonymous):

:) Some of them are funny. Help gotten and forgotten.